OpenStudy (anonymous):

If y=3x^50, then d^50y/dx^50=?

OpenStudy (amistre64):

hmm, the 50th derivative of a monomial ...

OpenStudy (anonymous):

Is it 3(50!)?

OpenStudy (amistre64):

maybe :) y^(0) = cx^n y^(1) = cnx^(n-1) y^(2) = cn(n-1)x^(n-2) y^(3) = cn(n-1)(n-2)x^(n-3) ... y^(k) = cn(n-1)(n-2)...(n-k+1)x^(n-k) seems right to me

OpenStudy (anonymous):

it would be 3(50!)